Let φ be the function on R^n defined by
φ(x) = e^(-|x|^(2)/2, x∈R
Compute (φ*φ)(x) for all x in R^n.
1
Expert's answer
2019-10-08T07:01:51-0400

"(\varphi*\varphi)(x)=\int\limits_{\mathbb R^n}\varphi(y)\cdot\varphi(x-y)dy=\int\limits_{\mathbb R^n}e^{-\frac{|y|^2}{2}}e^{-\frac{|x-y|^2}{2}}dy"

Let "y_0=(y_1,\ldots,y_{n-1})" and "x_0=(x_1,\ldots,x_{n-1})" , then "\int\limits_{\mathbb R^n}e^{-\frac{|y|^2}{2}}e^{-\frac{|x-y|^2}{2}}dy=\int\limits_{\mathbb R}\int\limits_{\mathbb R^{n-1}}e^{-\frac{|y_0|^2+y_n^2}{2}}e^{-\frac{|x_0-y_0|^2+(x_n-y_n)^2}{2}}dy_0dy_n="

"=\int\limits_{\mathbb R}\Bigl(\int\limits_{\mathbb R^{n-1}}e^{-\frac{|y_0|^2}{2}}e^{-\frac{|x_0-y_0|^2}{2}}dy_0\Bigr)e^{-\frac{y_n^2}{2}}e^{-\frac{(x_n-y_n)^2}{2}}dy_n="

"=\int\limits_{\mathbb R^{n-1}}e^{-\frac{|y_0|^2}{2}}e^{-\frac{|x_0-y_0|^2}{2}}dy_0\cdot\int\limits_{\mathbb R}e^{-\frac{y_n^2}{2}}e^{-\frac{(x_n-y_n)^2}{2}}dy_n"

Calculate "\int\limits_{\mathbb R}e^{-\frac{y_n^2}{2}}e^{-\frac{(x_n-y_n)^2}{2}}dy_n" :

"\int\limits_{\mathbb R}e^{-\frac{y_n^2}{2}}e^{-\frac{(x_n-y_n)^2}{2}}dy_n=\int\limits_{\mathbb R}e^{-\frac{2y_n^2+x_n^2-2x_ny_n}{2}}dy_n="

"=\int\limits_{\mathbb R}e^{-\frac{\Bigl(y_n\sqrt{2}-\frac{x_n}{\sqrt{2}}\Bigr)^2+\frac{x_n^2}{2}}{2}}dy_n"

Let "z_n=y_n\sqrt{2}-\frac{x_n}{\sqrt{2}}" , then "dy_n=\frac{1}{\sqrt{2}}dz_n" and

"\int\limits_{\mathbb R}e^{-\frac{\Bigl(y_n\sqrt{2}-\frac{x_n}{\sqrt{2}}\Bigr)^2+\frac{x_n^2}{2}}{2}}dy_n=\int\limits_{\mathbb R}e^{-\frac{z_n^2+\frac{x_n^2}{2}}{2}}\cdot\frac{1}{\sqrt{2}}dz_n="

"=\frac{1}{\sqrt{2}}e^{-\frac{x_n^2}{4}}\int\limits_{\mathbb R}e^{-\frac{z_n^2}{2}}dz_n"

We know that the probability density function of a normal disribution "N(0,1)" is "\rho(u)=\frac{1}{\sqrt{2\pi}}e^{-\frac{u^2}{2}}" , so "\int\limits_{\mathbb R}\frac{1}{\sqrt{2\pi}}e^{-\frac{u^2}{2}}du=1" , that is "\int\limits_{\mathbb R}e^{-\frac{u^2}{2}}du=\sqrt{2\pi}" . We obtain "\frac{1}{\sqrt{2}}e^{-\frac{x_n^2}{4}}\int\limits_{\mathbb R}e^{-\frac{z_n^2}{2}}dz_n=\frac{1}{\sqrt{2}}e^{-\frac{x_n^2}{4}}\sqrt{2\pi}=e^{-\frac{x_n^2}{4}}\sqrt{\pi}"

So "\int\limits_{\mathbb R^n}e^{-\frac{|y|^2}{2}}e^{-\frac{|x-y|^2}{2}}dy=e^{-\frac{x_n^2}{4}}\sqrt{\pi}\int\limits_{\mathbb R^{n-1}}e^{-\frac{|y_0|^2}{2}}e^{-\frac{|x_0-y_0|^2}{2}}dy"

Therefore we can prove by induction on "n" that "(\varphi*\varphi)(x)=\pi^{\frac{n}{2}}e^{-\frac{|x|^2}{4}}"

It is true for "n=1" : we proved "(\varphi*\varphi)(x)=\int\limits_{\mathbb R}e^{-\frac{y^2}{2}}e^{-\frac{(x-y)^2}{2}}dy=e^{-\frac{x^2}{4}}\sqrt{\pi}"

Suppose that it is true for "n=k", where "k\ge 1" . Then "\int\limits_{\mathbb R^{k+1}}e^{-\frac{|y|^2}{2}}e^{-\frac{|x-y|^2}{2}}dy=e^{-\frac{x_{k+1}^2}{4}}\sqrt{\pi}\int\limits_{\mathbb R^k}e^{-\frac{|y_0|^2}{2}}e^{-\frac{|x_0-y_0|^2}{2}}dy" , where "y_0=(y_1,\ldots,y_k)" and "x_0=(x_1,\ldots,x_k)" .

By induction hypohesis "\int\limits_{\mathbb R^k}e^{-\frac{|y_0|^2}{2}}e^{-\frac{|x_0-y_0|^2}{2}}dy=\pi^{\frac{k}{2}}e^{-\frac{|x_0|^2}{4}}" , so "e^{-\frac{x_{k+1}^2}{4}}\sqrt{\pi}\int\limits_{\mathbb R^k}e^{-\frac{|y_0|^2}{2}}e^{-\frac{|x_0-y_0|^2}{2}}dy=e^{-\frac{x_{k+1}^2}{4}}\sqrt{\pi}\cdot\pi^{\frac{k}{2}}e^{-\frac{|x_0|^2}{4}}="

"=\pi^{\frac{k+1}{2}}e^{-\frac{|x|^2}{4}}" . We proved that "\int\limits_{\mathbb R^{k+1}}e^{-\frac{|y|^2}{2}}e^{-\frac{|x-y|^2}{2}}dy=\pi^{\frac{k+1}{2}}e^{-\frac{|x|^2}{4}}".

By principle of mathematical induction we obtain "(\varphi*\varphi)(x)=\int\limits_{\mathbb R^n}e^{-\frac{|y|^2}{2}}e^{-\frac{|x-y|^2}{2}}dy=\pi^{\frac{n}{2}}e^{-\frac{|x|^2}{4}}"

Answer: "(\varphi*\varphi)(x)=\pi^{\frac{n}{2}}e^{-\frac{|x|^2}{4}}"